• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lhuillier

    Log entry time 16:00:07 on June16,2004

    Entry number 125970

    keyword=Shift Summary

    Prodution data till 8H45.
    Q1 trip --> go to controlled access.
    Q1 recovered well after reset in hall A.
    Then the septum software crashed. Helium levels in left Q1 and left septum went very low. Techs had to work in hall A to operate manually the valves and refill the helium reservoir of the left septum magnet. The remote control was broken and the PLC unit has been replaced. 
    At the end of this shift hall A techs are still working on the configuration of the CPU of the PLC.
    Hall A has been in controlled access since 9am with a brief period of beam around 9:30 am.